OpenStudy (anonymous):

A crow is flying horizontally with a constant speed of 2.45 m/s when it releases a clam from its beak. The clam lands on the rocky beach 2.30 s later. Consider the moment just before the clam lands. (Neglect air resistance.) a)What is its horizontal component of velocity? (in m/s) (b) What is its vertical component of velocity? (in m/s)

OpenStudy (matt101):

Neglecting air resistance, the horizontal velocity will not have changed (i.e. it remains 2.45 m/s). The vertical velocity starts at 0 m/s and accelerates for 2.30 s (due to gravity). 9.8 = (x-0)/2.3 x = 22.54 Therefore your vertical velocity is 22.54 m/s.

OpenStudy (anonymous):

a) It would remain the same, because there is no force, and thus acceleration, in that direction. So, it would just be 2.45 m/s. b) Use\[ v=v_{0}+at\] Because \[ v_{0}\]=0 in the y direction, it is just v=at It said that t=2.3 s, and a=9.8. Plug that in... v=2.3*9.8=22.54 m/s